OpenStudy (anonymous):

Karen bakes cookies. She has observed that the number of cookies that turn out well in each batch varies with the size of the batch and is given by the function f(x) = 2x – 23, where x is the size of the batch (number of cookies in each batch). She never bakes fewer than 12 cookies in a batch. If the size of five batches is {20, 25, 26, 27, 34}, how many cookies turn out well in each batch?

OpenStudy (anonymous):

Just substitute the values in for the x. I'll do the first one for you and you do the others.\[f(20)=2(20)-23 \to f(20)=40-23 \to f(20)=17 \] So the size of the 20 batch is 17 cookies

OpenStudy (anonymous):

alright thanks

OpenStudy (anonymous):

No problem :)